OpenStudy (anonymous):

sqrt(x+64)-sqrt(x-16)=4 Solve. I can't figure this out for the life of me.

hartnn (hartnn):

1) isolate any one of the square roots 2) then square both sides 3) repeat 1) 4) repeat 2)

OpenStudy (anonymous):

I tried that and I just got 64 and it says its wrong

hartnn (hartnn):

\(x+64 = (\sqrt {x-16}+4)^2 \) expand right side

hartnn (hartnn):

(a+b)^2 = a^2+ 2ab +b^2 a= sqrt (x-16) b=4

OpenStudy (triciaal):